# Tiling spaces are covering spaces over irrational tori

### v1

@@ 2.2 Background on Diffeology
Diffeological spaces are a generalization of differential manifolds. One of its attributes is that it is possible to deal with pathological topological spaces such as quotients by dense subspaces, the case we are interested in.
@@ 2.2.1 Diffeological Spaces
In a similar manner as in topology, diffeologies can be ordered, and there is a minimum and a maximum for this order. The minimum is the the fine or discrete diffeology whose plots are all locally constant; and the maximum is the coarse or indiscrete diffeology and is defined as the diffeology that contain every parametrization.
Diffeological spaces and smooth maps between them form a category which we shall denote by $\bf{Diff}$. It is a well-behaved category as it is closed under taking quotients, subobjects, limits, colimits and spaces of functions $C^{\infty}(\star,\star)$, where $C^{\infty}$ denotes the set of smooth maps between spaces in the diffeological sense. Natural structures are given to them via the basic constructions pull-back and push-forward of diffeologies from a space $(X,\mathcal{D})$ and maps $f:Y\to X$ and $g:X\to Z$, respectively. The pull-back diffeology $f^{*}(\mathcal{D})$ along $f$ (resp. push-forward $g_{*}(\mathcal{D})$ through $g$) are defined as the coarsest diffeology on $Y$ (resp. the finest diffeology on $Z$) such that $f$ (resp. $g$) is smooth. The plots for these diffeologies are specified as follows:
@@
The monomorphisms of $\bf{Diff}$ are called inductions: injective maps $i$ where the diffeology of the domain coincides with the pull-back along $i$ of the diffeology of the range; and the epimorphisms subductions: surjective maps $p$ where the diffeology of the range coincides with the push-forward through $p$ of the diffeology of the domain. The isomorphisms of $\bf{Diff}$ are called diffeomorphisms, they are smooth invertible maps whose inverse function is again smooth.
@@
For a smooth surjection $\pi:X\to B$ between diffeological spaces $X,B$ it is possible to define the structure groupoid $K_{\pi}$ of the surjection $\pi$ as the groupoid whose objets are $\mathrm{Obj}(K_{\pi})=B$ and its arrows are $\mathrm{Mor}(K_{\pi})=\bigcup_{b,b^{\prime}\in B}C^{\infty}(X_{b},X_{b^{\prime}})$, where $X_{b}$ denotes the fiber $\pi^{-1}(b)$. See [21, Postulate 8.2].
@@
The diffeological structure over a space induces naturally a topology, the so-called $D$-topology, whose open sets $A\subset X$ are precisely the subsets whose pre-images $P^{-1}(A)$ under every plot $P$ are open subsets of some $\mathbb{R}^{n}$. The $D$-topology over a diffeological space is the coarsest topology such that every plot is continuous.
The existence of the $D$-topology allows us to talk about local diffeological properties, defined in $D$-neighborhoods. A map $f:X\to X^{\prime}$ between diffeological spaces is said to be locally smooth in $x\in X$ if there is a $D$-open subset $A$ containing $x$ such that the map $f|_{A}$ is smooth. That is, if for every plot $P$ of $X$ the parametrization $f\circ P|_{P^{-1}(A)}$ is a plot of $X^{\prime}$. The map $f:X\to X^{\prime}$ is smooth if and only if it is locally smooth in every point of $X$. See [21, Postulate 2.3].
@@ 2.2.2 Diffeological Fiber Bundles
Let $\pi:X\to B$ be a smooth surjection between diffeological spaces. Then $\pi$ is a diffeological fibration if and only if the characteristic map $\chi$ of the structure groupoid $K_{\pi}$ (see Eq. 5) is a subduction. In such case $K_{\pi}$ is called fibrant.
A morphism between surjections $\pi:X\to B$ and $\pi^{\prime}:X^{\prime}\to B^{\prime}$ is a pair of smooth maps $(\Phi,\phi)$ making commutative the following diagram.
@@
Whenever $\Phi$ and $\phi$ are diffeomorphisms we call $\pi$ and $\pi^{\prime}$ equivalent. In particular, $\pi$ is trivial with diffeological fiber $F$ if it is equivalent to the first factor projection $\mathrm{pr}_{1}:B\times F\to B$. And, we say that $\pi$ is locally trivial with fiber $F$ if there exists a $D$-open cover $\{U_{i}\}_{i\in\Lambda}$ of $B$, such that the restriction $\pi|_{U_{i}}:\pi^{-1}(U_{i})\to U_{i}$ is a trivial surjection with fiber $F$.
When $\pi:X\to B$ is a diffeological fibration the surjectivity of $\chi$ implies that the preimages $X_{b}=\pi^{-1}(b)$ are all diffeomorphic: they are the fibers $F=X_{b}$.
@@
[21, Postulate 8.9] Let $X$ and $B$ diffeological spaces. A smooth map $\pi:X\to B$ is a fibration if and only if there exists a diffeological spaces $F$ such that the pullback $P^{*}(X)$ of $\pi$ by any plot $P:U\to B$ of $B$ is locally trivial with fiber $F$, where
$P^{*}(X):=\{(u,x)\in U\times X|P(u)=\pi(x)\}.$
@@
[21, Postulate 8.11] Consider the projection $\pi:X\to X/G$. If $\mathcal{A}_{X}^{G}$ is an induction then $\pi$ is a diffeological fibration with fiber $G$ and the action of $G$ on $X$ is called principal. Moreover, any equivalent fiber bundle $\pi^{\prime}:X\to Q$ to $\pi$ is a $G$-principal fiber bundle or $G$-principal fibration.
@@ 2.2.3 The irrational torus
is a diffelogical space and it is diffeomorphic to $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ (see [18, Proposition 3]). The density of $\mathbb{Z}+\alpha\mathbb{Z}$ in $\mathbb{R}$ implies that $\pi_{\alpha}$ is locally trivial along the plots of $\mathbb{T}_{\alpha}$ (see Lemma 2) but it is not locally trivial.
For two irrational numbers $\alpha,\beta$ the following lemma dictates the conditions for the irrational tori $\mathbb{T}_{\alpha},\mathbb{T}_{\beta}$ to be diffeomorphic.
@@
    There is a non-trivial smooth function $f\in C^{\infty}(\mathbb{T}_{\alpha},\mathbb{T}_{\beta})$ if and only if there are integers $a,b,c,d$ such that
@@
    The diffeomorphism $f$ induces an equivalence between the $\mathbb{R}$-principal fibrations $\pi_{\alpha}$ and $\pi_{\beta}$.
@@
for $x\in U$ and $n+m\alpha$ fixed. The irrationality of $\beta,\alpha$ implies that $n^{\prime}$ and $m^{\prime}$ are unique. But $F$ is smooth, then its derivative with respect to $x$ at $x_{0}$ gives us
@@
Using Equation 9, we get that
@@
Considering $\mathbb{Z}+\alpha\mathbb{Z},\mathbb{Z}+\beta\mathbb{Z}$ as $\mathbb{Z}$ modules, the smooth map $F$ is a isomorphism if and only if $ad-cb=\pm 1$.
@@
In particular, the orbit of $L_{\alpha}$ under the action of the group $\mathrm{GL}(2,\mathbb{Z})$ is a family of irrational tori diffeomorphic to $\mathbb{T}_{\alpha}$. The group $C^{\infty}(\mathbb{T}_{\alpha},\mathbb{T}_{\alpha})$ is a subgroup of $\mathrm{GL}(2,\mathbb{Z})$ and it is different from the group $\{\mathrm{id},\mathrm{-id}\}$ if and only if $\alpha$ is a quadratic irrational.
The diffeological fiber bundles satisfy the exact sequence of homotopy (see [21, Postulate 8.21]), which in our case gives us the following exact sequence:
@@
The line $\mathbb{R}$ as diffeological space is connected and simple connected, then the fundamental groups of $\mathbb{T}^{2}$ and $\mathbb{T}_{\alpha}$ are isomorphic.
@@ 3 Diffeological Tiling Spaces
In the present section we introduce a structure of diffeological space for the tiling spaces $\Omega$. For now, we let $\Omega$ be any space of tilings of $\mathbb{R}^{d}$.
@@ 3.1 The Diffeological Space $\Omega$
Define a diffeology $\mathcal{D}$ for $\Omega$ as the generated by the family of parametrizations
=$\mathcal{F}=\{F_{T}:U\subset\mathbb{R}^{d}\to\Omega:T\in\Omega\text{ and }F_{T}(x)=T+x\}.$
By definition of generating family [21, Postulate 1.68], the plots of $\mathcal{D}$ are specified as the parametrizations $P:U\to\Omega$ such that for every $u\in U$ there exists an open neighborhood $V\subset U$ of $u$ such that either $P|_{V}$ is a constant parametrization or there exist an element $F_{T}:W\to\Omega$ of $\mathcal{F}$ and a smooth map $Q:V\to W$ such that $P|_{V}=F_{T}\circ Q$. As shown in the following diagram.
@@
Note that the dimension of this generating family, defined as the supremum of the dimensions of its domains, is $d$. And, as no family of dimension less than $d$ may generate this diffeology, we deduce that the dimension of $\Omega$ is $d$ [21, Postulate 1.78].
The following couple of results verify that the natural actions by translation and substitution over $\Omega$ are smooth.
Lemma 4.
Let $\mathbf{Diff}(\Omega)$ be the group of diffeomorphisms of $\Omega$ endowed with the functional diffeology. Then the translation action $\rho:\mathbb{R}^{d}\to\mathbf{Diff}(\Omega)$ of $\mathbb{R}^{d}$ over $\Omega$, given by $\rho(x)(T)=T+x$, is smooth.
@@
We must show that the evaluation map
$ev_{\rho}:\Omega\times\mathbb{R}^{d}\to\Omega\text{ \hskip 0.56905pt given by \hskip 0.56905pt }ev_{\rho}(T,x)=\rho(x)(T)=T+x$ (14)
is smooth, i.e. for every plot $P$ of $\Omega\times\mathbb{R}^{d}$, the map $ev_{\rho}\circ P$ is a plot of $\Omega$. The latter condition happens if and only if every element $u\in U$ has a neighbourhood $V\subset U$ and there is a smooth map $f:V\to W$ between euclidian domains and an element $F_{T}:W\to\Omega$ of the generating family $\mathcal{F}$ such that
$(ev_{\rho}\circ P)|_{V}=F_{T}\circ f,$
that is, $ev_{\rho}\circ P$ is locally a translation up to pre-composing with a smooth map.
Let $P:U\to\Omega\times\mathbb{R}^{d}$ a plot in the product diffeology. This means that $\pi_{1}\circ P$ is a plot of $\Omega$ and $\pi_{2}\circ P:=h$ is a $C^{\infty}$ map from $U$ to $\mathbb{R}^{d}$, where $\pi_{i}$ denotes the projection onto the $i$-th factor. But being $\pi_{1}\circ P$ a plot for $\Omega$ means that it is, locally in a neighborhood $V$, a translation up to pre-composing with a smooth map, say $g$.
Then for every $u\in V\subset U$ we have that
$P(u)=(\pi_{1}\circ P(u),\pi_{2}\circ P(u))=(F_{T}\circ g(u),h(u))=(T+g(u),h(u))$
and replacing $P(u)$ in Equation 14 we get that
$ev_{\rho}\circ P(u)=\rho(h(u))(T+g(u))=T+g(u)+h(u)=F_{T}\circ(g+h)(u),$
where $g+h$ is a smooth map because both summands are smooth. Thus $ev_{\rho}\circ P$ is a plot for $\Omega$. Hence $ev_{\rho}$ is smooth and then $\rho$ acts smoothly.
Note that there are at least two natural topologies over $\Omega$: the $D$-topology and the topology induced by the metric $d$.
Let $A\subset\Omega$ be an open subset, assume without loss of generality that $A=\mathcal{B}_{\varepsilon}(T)$ for some $T\in\Omega$. Let $T^{\prime}\in A$, pick an $\varepsilon^{\prime}>0$ such that all translates of $T^{\prime}$ within radius $\varepsilon^{\prime}$ remain inside $A$. Let $B=\{T^{\prime}+x:\|x\|<\varepsilon^{\prime}\}\subset A$, we claim it is $D$-open. Let $P:U\to\Omega$ be a plot. By definition of a generating family (see discussion below Definition 12), for every $u\in U$, there exist a neighborhood $V_{u}\subset U$ of $u$ and a smooth map $Q_{u}:V_{u}\to U$ such that $P|_{V_{u}}=F_{T}\circ Q_{u}$. Then
$P^{-1}(B)=\bigcup_{u\in U}(P|_{V_{u}})^{-1}(B)=\bigcup_{u\in U}(F_{T}\circ Q_{u})^{-1}(B)=\bigcup_{u\in U}(Q_{u}^{-1}(F_{T}^{-1}(B)).$
But $F_{T}^{-1}(B)$ coincides with the euclidean open ball $B_{\varepsilon^{\prime}}(0)$ and $Q_{u}$ is a smooth map between euclidean domains therefore it is continuos. Thus $P^{-1}(B)$ is open in some $\mathbb{R}^{n}$ and $B$ is $D$-open. Thus the $D$-topology is finer than the topology induced by the metric $d$.
@@
If $\omega$ is a primitive substitution rule for $\Omega$ then the map $\omega:\Omega\to\Omega$ is smooth.
@@
Let $P:U\to\Omega$ be a plot for $\Omega$. We must show that $\omega\circ P$ is also a plot. By definition, for every point $u\in U$ there is a neighborhood $V\subset U$, an element $F_{T}:W\to\Omega$ of the generating family of $\mathcal{D}$ and a smooth map $Q:V\to W$ such that $P|_{V}=F_{T}\circ Q$ (assuming $P$ non constant, as if it were the result would follow).
[…]
@@
$\omega\circ P(v)=\omega\circ F_{T}\circ Q(v)=\omega\circ F_{T}(x)=\omega(T+x)=\omega(T)+\lambda\cdot x=F_{\omega(T)}(\lambda\cdot x),$
where $x=Q(v)$ and $\lambda$ is the expansion factor of $\omega$. The map $F_{\omega(T)}$ is an element of $\mathcal{F}$ because $\omega(T)\in\Omega$. It follows that every $u$ in $U$ has a neighborhood $V$ such that $\omega\circ P|_{V}=F_{\omega(T)}\circ Q$. Then $\omega\circ P$ is a plot for $\Omega$. ∎
Recall that the topological space $\Omega$ is locally homeomorphic to a cartesian product of a ball $D$ times totally disconnected set $K$. Note that Lemma 4 asserts that it is possible to move smoothly within the factor $D$, while the previous proposition means that there is also a smooth dynamic within the factor $K$.
With regard to equivalence we note that diffeomorphism of tiling spaces map orbits onto orbits, and viceversa, an orbital equivalence between tiling spaces is a diffeomorphism.
Proposition 6.
Let $(\Omega,\rho)$ and $(\Omega^{\prime},\rho^{\prime})$ be two $d$-dimensional tiling spaces endowed with their respective $\mathbb{R}^{d}$-actions $\rho$ and $\rho^{\prime}$ by translation. Then, they are orbitally equivalent if and only if they are diffeomorphic.
Proof.
Let $\varphi:\Omega\to\Omega^{\prime}$ an orbital equivalence between tiling spaces and $P:U\to\Omega$ a plot. Up to pre-composing with a smooth map $f$, $P$ is locally a translation. This means that for every $u\in U$ there exists an open neighborhood $V\subset U$ of $u$, a tiling $T\in\Omega$ and a smooth map $f:V\to W$ such that $P|_{V}=F_{T}\circ f$. Hence
$\varphi\circ P(u)=\varphi\circ F_{T}\circ f(u)=\varphi(T+f(u))=\varphi(T)+y=F^{\prime}_{\varphi(T)}(y),$
for some $y\in\mathbb{R}^{d}$. The third equality is given by hypothesis and $F^{\prime}_{\varphi(T)}$ is an element of the generating family $\mathcal{F}^{\prime}$ for the diffeology of $\Omega^{\prime}$. Thus $\varphi\circ P$ is a plot of $\Omega^{\prime}$ and then $\varphi$ is smooth. The same argument works for proving the smoothness of $\varphi^{-1}$. Hence $\varphi$ is a diffeomorphism.
Conversely, let $\varphi:\Omega\to\Omega^{\prime}$ be a diffeomorphism. Let $T\in\Omega$ and $x\in\mathbb{R}^{d}$. Consider the path $\gamma:\mathbb{R}\to\Omega$ given by $\gamma(s)=(1-s)T+s(T+x)$. This path joins $T$ to $T+x$. Then, as $\varphi$ is smooth, $\varphi\circ\gamma$ is a path in $\Omega^{\prime}$ such that $\varphi\circ\gamma(0)=\varphi(T)$ and $\varphi\circ\gamma(1)=\varphi(T+x)$. Thus $\varphi(T)$ and $\varphi(T+x)$ lie in the same path component, so $\varphi(T+x)$ must be a translate of $\varphi(T)$. So $\varphi$ takes orbits in $\Omega$ onto orbits in $\Omega^{\prime}$.
3.2 The bundle $\hat{p}:\Omega\to\mathbb{T}_{\alpha}$
In the following we restrict to one-dimensional tilings. We will discuss on how to bring Sadun-Williams’ construction into the diffeological setting. Consider an unidimensional tiling space $\Omega$ of primitive substitution tilings with a rule $\omega$ on a two-tile set $\{0,1\}$. Recall from [2] that $\Omega$ is conjugate – and hence diffeomorphic following Proposition 6– to the inverse limit of branched curves $\Omega_{n}$

### v3

@@ 2.2 Background on Diffeology
Diffeology is a generalization of differential geometry that includes manifolds and orbifolds, among other spaces. One of its attributes is that it is possible to deal with pathological spaces such as quotients by dense subspaces – the case we are interested in.
@@ 2.2.1 Diffeological Spaces
In a similar manner as in topology, diffeologies can be ordered and there is a minimum and a maximum with respect to this order. The minimum is the the fine or discrete diffeology whose plots are all locally constant; and the maximum is the coarse or indiscrete diffeology and it is defined as the diffeology that contains every parametrization.
Diffeological spaces and smooth maps between them form a category which we shall denote by $\bf{Diff}$. It is a well-behaved category in the sense that it is closed under taking quotients, subobjects, limits, colimits and spaces of functions $C^{\infty}(\star,\star)$, where $C^{\infty}$ denotes the set of smooth maps between spaces in the diffeological sense. Natural structures are given to them via the basic constructions pull-back and push-forward of diffeologies from a space $(X,\mathcal{D})$ and maps $f:Y\to X$ and $g:X\to Z$, respectively. The pull-back diffeology $f^{*}(\mathcal{D})$ along $f$ (resp. push-forward $g_{*}(\mathcal{D})$ through $g$) is defined as the coarsest diffeology on $Y$ (resp. the finest diffeology on $Z$) such that $f$ (resp. $g$) is smooth. The plots for these diffeologies are specified as follows:
@@
The monomorphisms of $\bf{Diff}$ are called inductions: injective maps $i$ where the diffeology of the domain coincides with the pull-back along $i$ of the diffeology of the range; and the epimorphisms subductions: surjective maps $p$ where the diffeology of the range coincides with the push-forward through $p$ of the diffeology of the domain. The isomorphisms of $\bf{Diff}$ are called diffeomorphisms. They are smooth invertible maps whose inverse function is again smooth.
@@
For a smooth surjection $\pi:X\to B$ between diffeological spaces $X,B$ it is possible to define the structure groupoid $K_{\pi}$ of the surjection $\pi$ as the groupoid whose objects are $\mathrm{Obj}(K_{\pi})=B$ and its arrows are $\mathrm{Mor}(K_{\pi})=\bigcup_{b,b^{\prime}\in B}C^{\infty}(X_{b},X_{b^{\prime}})$, where $X_{b}$ denotes the fiber $\pi^{-1}(b)$. See [19, Postulate 8.2].
@@
The diffeological structure over a space naturally induces a topology, the so-called $D$-topology, whose open sets $A\subset X$ are precisely the subsets whose pre-images $P^{-1}(A)$ under every plot $P$ are open subsets of some $\mathbb{R}^{n}$. The $D$-topology over a diffeological space is the coarsest topology such that every plot is continuous.
The existence of the $D$-topology allows us to talk about local diffeological properties, defined in $D$-neighborhoods. A map $f:X\to X^{\prime}$ between diffeological spaces is said to be locally smooth in $x\in X$ if there is a $D$-open subset $A$ containing $x$ such that the map $f|_{A}$ is smooth. That is, if for every plot $P$ of $X$ the parametrization $f\circ P|_{P^{-1}(A)}$ is a plot of $X^{\prime}$. The map $f:X\to X^{\prime}$ is smooth if and only if it is locally smooth in every point of $X$. See [19, Postulate 2.3].
@@ 2.2.2 Diffeological Fiber Bundles
Let $\pi:X\to B$ be a smooth surjection between diffeological spaces. Then $\pi$ is a diffeological fibration if and only if the characteristic map $\chi$ of the structure groupoid $K_{\pi}$ (see Eq. 5) is a subduction. In such case, $K_{\pi}$ is called fibrating.
A morphism between surjections $\pi:X\to B$ and $\pi^{\prime}:X^{\prime}\to B^{\prime}$ is a pair of smooth maps $(\Phi,\phi)$ making the following diagram commute.
@@
Whenever $\Phi$ and $\phi$ are diffeomorphisms we say that $\pi$ and $\pi^{\prime}$ are equivalent. In particular, $\pi$ is trivial with diffeological fiber $F$ if it is equivalent to the projection $\mathrm{pr}_{1}:B\times F\to B$. We say that $\pi$ is locally trivial with fiber $F$ if there exists a $D$-open cover $\{U_{i}\}_{i\in\Lambda}$ of $B$ such that the restrictions $\pi|_{U_{i}}:\pi^{-1}(U_{i})\to U_{i}$ are trivial surjections with fiber $F$.
When $\pi:X\to B$ is a diffeological fibration the surjectivity of $\chi$ (5) implies that the preimages $X_{b}=\pi^{-1}(b)$ are all diffeomorphic: they are the fibers $F=X_{b}$.
@@
[19, Postulate 8.9] Let $X$ and $B$ diffeological spaces. A smooth map $\pi:X\to B$ is a fibration if and only if there exists a diffeological space $F$ such that the pullback $P^{*}(X)$ of $\pi$ along any plot $P:U\to B$ of $B$ is locally trivial with fiber $F$, where
$P^{*}(X):=\{(u,x)\in U\times X:P(u)=\pi(x)\}.$
@@
[19, Postulate 8.11] Consider the projection $\pi:X\to X/G$. If $\mathcal{A}_{X}^{G}$ is an induction then $\pi$ is a diffeological fibration with fiber $G$ and the action of $G$ on $X$ is called principal. Moreover, any equivalent fiber bundle $\pi^{\prime}:X\to Q$ to $\pi$ is a $G$-principal fiber bundle or $G$-principal fibration.
@@ 2.2.3 The irrational torus
is a diffeological space and it is diffeomorphic to $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ (see [18, Proposition 3]). The density of $\mathbb{Z}+\alpha\mathbb{Z}$ in $\mathbb{R}$ implies that $\pi_{\alpha}$ is locally trivial along the plots of $\mathbb{T}_{\alpha}$ (see Lemma 2) but it is not locally trivial.
For two irrational numbers $\alpha,\beta$, the following lemma from [18] dictates the conditions for the irrational tori $\mathbb{T}_{\alpha},\mathbb{T}_{\beta}$ to be diffeomorphic.
@@
    There is a non-trivial smooth map $f\in C^{\infty}(\mathbb{T}_{\alpha},\mathbb{T}_{\beta})$ if and only if there are integers $a,b,c,d$ such that
@@
    If (iii) hold, the diffeomorphism $f$ induces an equivalence between the $\mathbb{R}$-principal fibrations $\pi_{\alpha}$ and $\pi_{\beta}$.
@@
for $x\in U$ and $n+m\alpha$ fixed. The irrationality of $\alpha,\beta$ implies that $n^{\prime}$ and $m^{\prime}$ are unique. But $F$ is smooth, then its derivative with respect to $x$ at $x_{0}$ gives us
@@
Using Equation 9 we get that
@@
Considering $\mathbb{Z}+\alpha\mathbb{Z},\mathbb{Z}+\beta\mathbb{Z}$ as $\mathbb{Z}$ modules, the smooth map $F$ is an isomorphism if and only if $ad-cb=\pm 1$.
@@
In particular, the orbit of $L_{\alpha}$ under the action of the group $\mathrm{GL}(2,\mathbb{Z})$ is a family of irrational tori diffeomorphic to $\mathbb{T}_{\alpha}$. The group of connected components of $C^{\infty}(\mathbb{T}_{\alpha},\mathbb{T}_{\alpha})$ is a subgroup of $\mathrm{GL}(2,\mathbb{Z})$ and it is different from the group $\{\mathrm{id},\mathrm{-id}\}$ if and only if $\alpha$ is a quadratic irrational (see Proposition 3.3 in [21]).
Another property worth mentioning is that the exact sequence of homotopy for fiber bundles also holds in the diffeological context, (see [19, Postulate 8.21]). In our case, it gives us the following exact sequence:
@@
The diffeological space $\mathbb{R}$ is connected and simply connected, then the fundamental groups of $\mathbb{T}^{2}$ and $\mathbb{T}_{\alpha}$ are isomorphic.
@@ 3 Diffeological Tiling Spaces
In the present section we introduce a diffeological structure for the tiling spaces $\Omega$. For now, we let $\Omega$ be any space of tilings of $\mathbb{R}^{d}$.
@@ 3.1 The Diffeological Space $\Omega$
Define a diffeology $\mathcal{D}$ for $\Omega$ as the finest diffeology such that the translation action of $\mathbb{R}^{d}\times\Omega\to\Omega$ given by $(x,T)\mapsto T+x$ is smooth.
Observe that, for any diffeology $\mathcal{D}^{\prime}$ on $\Omega$ such that the action of $\mathbb{R}^{d}$ by translation is smooth, the orbit maps $\hat{T}:\mathbb{R}^{d}\to\Omega$ given by $\hat{T}(x)=T+x$ are smooth for every $T$. This means that any such diffeology $\mathcal{D}^{\prime}$ contains every orbit map $\hat{T}:\mathbb{R}^{d}\to\Omega$ as a plot, and thus the diffeology generated by the family of orbit maps is contained in $\mathcal{D}^{\prime}$. Let
=$\mathcal{F}=\{\hat{T}:\mathbb{R}^{d}\to\Omega:T\in\Omega\text{ and }\hat{T}(x)=T+x\}.$ (14)
and $\mathcal{D}_{1}=\langle F\rangle$. We have that the diffeology $\mathcal{D}$ from Definition 12 contains $\mathcal{D}_{1}$. But $\mathcal{D}$ is minimal by definition of generating family. Thus $\mathcal{D}=\mathcal{D}_{1}$.
By definition of generating family [19, Postulate 1.68], the plots of $\mathcal{D}=\langle F\rangle$ are specified as the parametrizations $P:U\to\Omega$ such that for every $u\in U$ there exists an open neighborhood $V\subset U$ of $u$ such that either $P|_{V}$ is a constant parametrization or there exists an element $\hat{T}:\mathbb{R}^{d}\to\Omega$ of $\mathcal{F}$ and a smooth map $Q:V\to\mathbb{R}^{d}$ such that $P|_{V}=\hat{T}\circ Q$. As shown in the following diagram.
@@
With regard to equivalence we note that diffeomorphisms of tiling spaces map orbits onto orbits, and viceversa, an orbit equivalence between tiling spaces is a diffeomorphism.
Proposition 5.
Let $(\Omega,\rho)$ and $(\Omega^{\prime},\rho^{\prime})$ be two $d$-dimensional tiling spaces endowed with their respective $\mathbb{R}^{d}$-actions $\rho$ and $\rho^{\prime}$ by translation. Then, they are orbital equivalent if and only if they are diffeomorphic.
@@
Let $\varphi:\Omega\to\Omega^{\prime}$ an orbital equivalence between tiling spaces and $P:U\to\Omega$ a plot. Up to pre-composing with a smooth map $f$, $P$ is locally a translation. This means that for every $u\in U$ there exists an open neighborhood $V\subset U$ of $u$, a tiling $T\in\Omega$ and a smooth map $f:V\to\mathbb{R}^{d}$ such that $P|_{V}=\hat{T}\circ f$. Hence
$\varphi\circ P(u)=\varphi\circ\hat{T}\circ f(u)=\varphi(T+f(u))=\varphi(T)+y=\hat{\varphi(T)}(y),$
for some $y\in\mathbb{R}^{d}$. The third equality is given by hypothesis and $\hat{\varphi(T)}$ is an element of the generating family $\mathcal{F}^{\prime}$ for the diffeology of $\Omega^{\prime}$. Thus $\varphi\circ P$ is a plot of $\Omega^{\prime}$ and then $\varphi$ is smooth. The same argument works for proving the smoothness of $\varphi^{-1}$. Hence $\varphi$ is a diffeomorphism.
Conversely, let $\varphi:\Omega\to\Omega^{\prime}$ be a diffeomorphism. Let $T\in\Omega$ and $x\in\mathbb{R}^{d}$. Consider the path $\gamma:\mathbb{R}\to\Omega$ given by $\gamma(s)=(1-s)T+s(T+x)$. This path joins $T$ to $T+x$. Then, as $\varphi$ is smooth, $\varphi\circ\gamma$ is a path in $\Omega^{\prime}$ such that $\varphi\circ\gamma(0)=\varphi(T)$ and $\varphi\circ\gamma(1)=\varphi(T+x)$. Thus $\varphi(T)$ and $\varphi(T+x)$ lie in the same path component of $\Omega^{\prime}$, so $\varphi(T+x)$ must be a translate of $\varphi(T)$. Hence $\varphi$ maps orbits in $\Omega$ onto orbits in $\Omega^{\prime}$.
@@
If $\omega$ is a primitive substitution rule for $\Omega$ then the map $\omega:\Omega\to\Omega$ is smooth. Moreover, if $\Omega$ contains at least one non-periodic tiling then $\omega$ is a diffeomorphism.
@@
Let $P:U\to\Omega$ be a plot for $\Omega$. We must show that $\omega\circ P$ is also a plot. By definition, for every point $u\in U$ there is a neighborhood $V\subset U$, an element $\hat{T}:\mathbb{R}^{d}\to\Omega$ of the generating family of the diffeology $\mathcal{D}$ and a smooth map $Q:V\to W$ such that $P|_{V}=\hat{T}\circ Q$ (assuming $P$ non constant, as if it were the result would follow).
${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P}$${\Omega\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\omega}$${\Omega}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{Q}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\hat{T}}$$\scriptstyle{\hat{\omega(T)}}$
@@
$\omega\circ P(v)=\omega\circ\hat{T}\circ Q(v)=\omega\circ\hat{T}(x)=\omega(T+x)=\omega(T)+\lambda\cdot x=\hat{\omega(T)}(\lambda\cdot x),$
where $x=Q(v)$ and $\lambda$ is the expansion factor of $\omega$. The map $\hat{\omega(T)}$ is an element of $\mathcal{F}$ because $\omega(T)\in\Omega$. It follows that every $u$ in $U$ has a neighborhood $V$ such that $\omega\circ P|_{V}=\hat{\omega(T)}\circ Q$. Then $\omega\circ P$ is a plot for $\Omega$ and $\omega$ is smooth.
The fact that $\omega$ is invertible whenever $\Omega$ contains a non-periodic tiling follows from Theorem 1.3 in [35], and the smoothness of $\omega^{-1}$ is verified in the same way as before. ∎
Recall that the topological space $\Omega$ is locally homeomorphic to a cartesian product of a ball $D$ times totally disconnected set $K$. Note that by the definition of $\mathcal{D}$ we know that it is possible to move smoothly within the factor $D$, while the previous proposition means that there is also a smooth dynamic within the factor $K$.
Remark 3.
Note that there are at least two natural topologies over $\Omega$: the $D$-topology and the topology induced by the metric $d$. It is not difficult to prove that the former is finer than the latter.
3.2 The bundle $\tilde{p}:\Omega\to\mathbb{T}_{\alpha}$
In the current section will discuss on how to bring Sadun-Williams’ (Theorem 3) construction into the diffeological setting. For simplicity, we consider first an unidimensional tiling space $\Omega$. Recall from Theorem 2 that $\Omega$ is conjugate –and hence diffeomorphic following Proposition 5– to the inverse limit of branched curves $\Omega_{n}$